OpenStudy (togundeji):

3. The ratio of the width to the length in rectangle ABCD is 4 colon 5 . If the area of the rectangle is 2,000 square meters, then which is the length of the rectangle? (1 point) 9 meters 10 meters 40 meters 50 meters

OpenStudy (togundeji):

@MadaraUchiha @ellamoyseyuk @Mjudson

OpenStudy (togundeji):

@winsteria @weirdjanel @quin100 @q@Elsa213 @EricaEllis2016 @EricaMarie @ellamoyseyuk @phi @paki @Photon336 @priyar @Peace420 @peachescherry14 @dancergirl45 @d@DallasCowboys88 @dengeki_daisy @dumbcow @samanthagreer @sissyedgar @Savannah_Noelle @zeus4263

OpenStudy (zeus4263):

why was i brought into this?

OpenStudy (mjdennis):

What have you tried so far?

OpenStudy (anonymous):

\[\frac{w}{l} = \frac{4x}{5x}\] \[A = w*l = 4x *5x = 20x^2\] \[20x^2 = 2000\] \[x^2 = 100\] \[x = 10\] what is the length?

OpenStudy (togundeji):

@zeus4263 I just hoped that you would be able to help

OpenStudy (togundeji):

@dumbcow Would it be 10

OpenStudy (togundeji):

@LunyMoony can you help?

OpenStudy (lunymoony):

Yes, it would be 10

OpenStudy (togundeji):

Thank you! You are a life saver! @LunyMoony